OpenStudy (anonymous):

An acute angle θ is in a right triangle with sin θ =6/7 . What is the value of cot θ? Is the answer 6/sqrt13?

OpenStudy (anonymous):

\[\cot \theta=\frac{ \cos \theta }{ \sin \theta }\]

OpenStudy (anonymous):



OpenStudy (anonymous):

cotx= 1/tanx= cosx/sinx

OpenStudy (anonymous):

the answer you got, 6/square route 13 is the tangent

OpenStudy (anonymous):

The reciprocal of your answer is correct answer.

OpenStudy (anonymous):

oh. im an idiot. sqrt13/6?

OpenStudy (anonymous):

@Caolco